2x-3y=6 graph the linear equation by first finding the intercepts
Answer by nerdybill(7384)

2x-3y=6 graph the linear equation by first finding the intercepts
.
To find the x-intercept (this is where the line crosses the x-axis), simply set y=0 and solve for x:
2x-3y=6
2x-3(0)=6
2x=6
x = 3
Therefore, the x-intercept is at (3,0)
.
To find the y-intercept (this is where the line crosses the y-axis), simply set x=0 and solve for y:
2x-3y=6
2(0)-3y=6
-3y=6
y= -2
Therefore, the y-intercept is at (0,-2)
.
To graph:
Using graph paper, plot the two points:
(3,0)
(0,-2)
Draw a solid line through those two points -- that's all there is to it.